# MCFEND: A Multi-source Benchmark Dataset for Chinese Fake News Detection

### v1

@@ 1. Introduction
To bridge this gap, we constructed the first Multi-source benchmark dataset for Chinese FakE News Detection (MCFEND), which contains 23,974 pieces of authoritatively verified Chinese news from 14 fact-checking agencies covering numerous news sources. [^3] The Weibo Community Management Center, Weibo’s official fact-checking agency, exclusively examines news sourced from Weibo, whereas other fact-checking agencies cover various news sources. Note that the inclusion of additional fact-checking agencies can potentially enhance the variety of news sources considered. Please refer to Table 2 for the full list of included fact-checking agencies. These fact-checking agencies are further divided into three distinct groups. The first group encompasses nine Chinese fact-checking agencies. That means the collected news has been verified by experts as active and authoritative. The second group corresponds to three existing annotated English fake news detection datasets. Specifically, for an English news piece paired with its corresponding authenticity label from an existing English fake news dataset, we employ the state-of-the-art cross-lingual identical news retrieval system to collect its Chinese equivalent while retaining its original label. The third group consists of Weibo’s official fact-checking agency exclusively, i.e., the Weibo Community Management Center. For this group of news, we utilized news data from the Weibo-21 dataset (Nan et al., 2021). Furthermore, we conducted comprehensive evaluations on six established baseline models for Chinese fake news detection, including state-of-the-art methods, under cross-source and multi-source scenarios on our MCFEND dataset. The experimental results characterize the challenge of accurately spotting fake news from different sources that the dataset presents.
Our contributions are summarized as follows: First, We constructed the initial multi-source Chinese fake news detection dataset MCFEND, which comprises multi-modal content and social context of 23,974 real-world Chinese news pieces collected from 14 authoritative fact-checking agencies in three distinct groups. Additionally, to the best of our knowledge, MCFEND is the largest open-sourced Chinese fake news detection dataset, being at least 2.63 times larger than any existing ones. The dataset aims to benchmark the evaluations of Chinese fake news detection methods in real-world scenarios, where news originates from diverse sources, and to encourage further research in this field. Second, we conducted comprehensive cross-source, multi-source, and unseen source evaluations on six established baseline models for Chinese fake news detection, including state-of-the-art methods. Our experimental results reveal that the models trained on existing datasets are not applicable in real-world scenarios. Incorporating multi-source data is necessary, which can enhance the models’ robustness substantially.
@@ 2. Preliminaries and Related work
Existing datasets for Chinese fake news detection rely heavily on Weibo. Different from them, we constructed the pioneering multi-source Chinese fake news detection dataset, termed MCFEND, which contains 23,974 real-world Chinese news pieces collected from multiple sources across three distinct categories. Table 1 compares the Chinese fake news detection datasets.
@@ 3.1. Overview
The MCFEND dataset contains news verified by 14 fact-checking agencies from a wide range of sources, such as messaging apps, social platforms, and traditional news outlets. As mentioned above, these agencies are categorized into three groups (see Table 2. The websites for the 14 fact-checking agencies are detailed in Appendix A. The first group includes nine Chinese fact-checking agencies identified as active by Duke Reporters [^7] https://reporterslab.org/fact-checking/, along with four Chinese fact-checking agencies manually verified by experts as active and authoritative. The second group corresponds to four English fact-checking agencies, including Politifact, Gossipcop, BS Detector, and FakeNewsCorpus. This group contains the Chinese counterparts of the English news fact-checked by the aforementioned agencies, which are collected through a carefully designed cross-lingual identical news retrieval method. An in-depth description of the method can be found in Sec. 3.2.2. Group 3 exclusively covers the Weibo Community Management Center, from which news data was directly sourced from Weibo-21 dataset (Nan et al., 2021).
@@
The overall MCFEND dataset contains 23,974 news pieces, including 8,144 sourced from the nine fact-checking agencies in Group 1, 6,702 related to the four English fact-checking agencies in Group 2, and 9,128 obtained from the Weibo Community Management Center in Group 3. Similar to existing dataset construction (Shu et al., 2020; Nan et al., 2021), we collected the following information for each piece of news in any group: (1) Multi-modal news content, including text, images, and metadata, e.g., timestamps; (2) Multi-modal social context, including posts, comments, emojis, user profiles, and other metadata, e.g., like counts of comments. Table 3 presents the detailed statistics of the MCFEND dataset.
@@ 3.2.1. Group 1: Fact-checking Agencies Data Crawling
To maximize our coverage of news sources, we conducted web crawling to collect data from all active Chinese fact-checking agencies, encompassing the five Chinese fact-checking agencies identified as active by Duke Reporters, in addition to nine other Chinese fact-checking agencies that were manually verified as active and authoritative. In the case where the labels on some fact-checking agencies, e.g., AFP Fact Check Asia and Factcheck Lab, are presented in the form of images, we utilized an optical character recognition method called Tesseract-OCR to retrieve such labels. [^8] https://github.com/tesseract-ocr/tesseract
@@ Appendix A The Fact-checking Agencies
China Internet Joint Rumor Refuting Platform
https://www.piyao.org.cn/
Tencent Jiaozhen
https://vp.fact.qq.com/
China Daily Factcheck
https://www.chinadaily.com.cn/china/factcheck/
Taiwan FactCheck Center
https://tfc-taiwan.org.tw/
MyGoPen
https://www.mygopen.com/
HKBU Factcheck
https://factcheck.hkbu.edu.hk/home/
HKU Annie Lab
https://annielab.org/
AFP Fact Check Asia
https://factcheck.afp.com/afp-asia/
Factcheck Lab
https://www.factchecklab.org/
Politifact
https://www.politifact.com/
Gossipcop
https://www.gossipcop.com/
BS Detector
https://github.com/selfagency/bs-detector
FakeNewsCorpus
https://github.com/architapathak/FakeNewsCorpus
Weibo Community Management Center
https://service.account.weibo.com
@@ Appendix C Experimental Setup
BERT
learning rate: 2e-5;
@@
RoBERTa
learning rate: 2e-5;
@@
CLIP
learning rate: 2e-5; project_dim: 512.
CAFE
learning rate: 1e-3;
@@
dEFEND
learning rate: 2e-5;
@@
BERT-EMO
learning rate: 1e-3;

### v2

@@ 1. Introduction
To bridge this gap, we constructed the first Multi-source benchmark dataset for Chinese FakE News Detection (MCFEND), which contains 23,789 pieces of authoritatively verified Chinese news from 14 fact-checking agencies covering numerous news sources. [^3] The Weibo Community Management Center, Weibo’s official fact-checking agency, exclusively examines news sourced from Weibo, whereas other fact-checking agencies cover various news sources. Note that the inclusion of additional fact-checking agencies can potentially enhance the variety of news sources considered. Please refer to Table 2 for the full list of included fact-checking agencies. These fact-checking agencies are further divided into three distinct groups. The first group encompasses nine Chinese fact-checking agencies. That means the collected news has been verified by experts as active and authoritative. The second group corresponds to three existing annotated English fake news detection datasets. Specifically, for an English news piece paired with its corresponding authenticity label from an existing English fake news dataset, we employ the state-of-the-art cross-lingual identical news retrieval system to collect its Chinese equivalent while retaining its original label. The third group consists of Weibo’s official fact-checking agency exclusively, i.e., the Weibo Community Management Center. For this group of news, we utilized news data from the Weibo-21 dataset (Nan et al., 2021). Furthermore, we conducted comprehensive evaluations on six established baseline models for Chinese fake news detection, including state-of-the-art methods, under cross-source and multi-source scenarios on our MCFEND dataset. The experimental results characterize the challenge of accurately spotting fake news from different sources that the dataset presents.
Our contributions are summarized as follows: First, we constructed the initial multi-source Chinese fake news detection dataset MCFEND, which comprises multi-modal content and social context of 23,789 real-world Chinese news pieces collected from 14 authoritative fact-checking agencies in three distinct groups. To the best of our knowledge, MCFEND is the largest open-sourced Chinese fake news detection dataset, being at least 2.63 times larger than any existing ones. The dataset aims to benchmark the evaluations of Chinese fake news detection methods in real-world scenarios, where news originates from diverse sources, and to encourage further research in this field. Second, we conducted comprehensive cross-source, multi-source, and unseen source evaluations on six established baseline models for Chinese fake news detection, including state-of-the-art methods. Our experimental results reveal that the models trained on existing datasets are not applicable in real-world scenarios. Incorporating multi-source data is necessary, which can enhance the models’ robustness substantially.
@@ 2. Preliminaries and Related work
Existing datasets for Chinese fake news detection rely heavily on Weibo. Different from them, we constructed the pioneering multi-source Chinese fake news detection dataset, termed MCFEND, which contains 23,789 real-world Chinese news pieces collected from multiple sources across three distinct categories. Table 1 compares the Chinese fake news detection datasets.
@@ 3.1. Overview
The MCFEND dataset contains news verified by 14 fact-checking agencies from a wide range of sources, such as messaging apps, social platforms, and traditional news outlets. As mentioned above, these agencies are categorized into three groups (see Table 2. The websites for the 14 fact-checking agencies are detailed in Appendix A. The first group includes five Chinese fact-checking agencies identified as active by Duke Reporters [^7] https://reporterslab.org/fact-checking/, along with four other Chinese fact-checking agencies manually verified by experts as active and authoritative. The second group corresponds to four English fact-checking agencies, including Politifact, Gossipcop, BS Detector, and FakeNewsCorpus. This group contains the Chinese counterparts of the English news fact-checked by the aforementioned agencies, which are collected through a carefully designed cross-lingual identical news retrieval method. An in-depth description of the method can be found in Sec. 3.2.2. Group 3 exclusively covers the Weibo Community Management Center, from which news data was directly sourced from Weibo-21 dataset (Nan et al., 2021).
@@
The overall MCFEND dataset contains 23,789 news pieces, including 7,959 sourced from the nine fact-checking agencies in Group 1, 6,702 related to the four English fact-checking agencies in Group 2, and 9,128 obtained from the Weibo Community Management Center in Group 3. Similar to existing dataset construction (Shu et al., 2020; Nan et al., 2021), we collected the following information for each piece of news in any group: (1) Multi-modal news content, including text, images, and metadata, e.g., timestamps; (2) Multi-modal social context, including posts, comments, emojis, user profiles, and other metadata, e.g., like counts of comments. Table 3 presents the detailed statistics of the MCFEND dataset.
@@ 3.2.1. Group 1: Fact-checking Agencies Data Crawling
To maximize our coverage of news sources, we conducted web crawling to collect data from all active Chinese fact-checking agencies, encompassing the five Chinese fact-checking agencies identified as active by Duke Reporters, in addition to four other Chinese fact-checking agencies that were manually verified as active and authoritative. In the case where the labels on some fact-checking agencies, e.g., AFP Fact Check Asia and Factcheck Lab, are presented in the form of images, we utilized an optical character recognition method called Tesseract-OCR to retrieve such labels. [^8] https://github.com/tesseract-ocr/tesseract
@@ Appendix A The Fact-checking Agencies
China Internet Joint Rumor Refuting Platform https://www.piyao.org.cn/
Tencent Jiaozhen https://vp.fact.qq.com/
China Daily Factcheck https://www.chinadaily.com.cn/china/factcheck/
Taiwan FactCheck Center https://tfc-taiwan.org.tw/
MyGoPen https://www.mygopen.com/
HKBU Factcheck https://factcheck.hkbu.edu.hk/home/
HKU Annie Lab https://annielab.org/
AFP Fact Check Asia https://factcheck.afp.com/afp-asia/
Factcheck Lab https://www.factchecklab.org/
Politifact https://www.politifact.com/
Gossipcop https://www.gossipcop.com/
BS Detector https://github.com/selfagency/bs-detector
FakeNewsCorpus https://github.com/architapathak/FakeNewsCorpus
Weibo Community Management Center https://service.account.weibo.com
@@ Appendix C Experimental Setup
BERT learning rate: 2e-5;
@@
RoBERTa learning rate: 2e-5;
@@
CLIP learning rate: 2e-5; project_dim: 512.
CAFE learning rate: 1e-3;
@@
dEFEND learning rate: 2e-5;
@@
BERT-EMO learning rate: 1e-3;